OpenStudy (anonymous):

Will Fan and medal 5. Jessica purchased a DVD that was on sale for 12% off. The sales tax in her county is 3%. Let y represent the original price of the DVD. Write an expression that can be used to determine the final cost of the DVD. (5 points)

OpenStudy (anonymous):

6. In Brock's class of 25 students, 15 students report that math is their favorite class and 40% of the students report that science is their favorite class. Show your work to explain how to determine whether more students prefer math or science. (5 points)

imqwerty (imqwerty):

\[P \frac{ 88 }{ 100 }\left[ 1+\frac{ 103 }{ 100} \right]\] where P is original price

OpenStudy (anonymous):

@imqwerty uhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh what lol